A Childs grandparents purchase a $16,000 bond that matures in 16 years to be used for her education. The bond pays 6% interest compound semiannually. How much will the bond be worth at maturity?
Answer by checkley71(8403)

I=P(1+R/N)^TN
I=16,000(1+.06/2)^16*2
I=16,000(1.03)^32
I=16,000*2.575
I=41,200
16,000+41,200=$57,200 IS THE VALUE OF THE BOND AFTER 16 YEARS COMPOUNDED SEMIANNUALLY.